PC is twice the length of a B in CD is twice the length of BC if a G is 140 inches long how many inches long is EF


1
Expert's answer
2021-02-09T03:29:53-0500

Let us try to bring more order into the condition to see how such problems can be solved:

PC is twice the length of AB in CD. CD is twice the length of BC. If AC is 140 inches long, how many inches long is BP?


Solution

1) Draw the situation and the known values:



What we need to find is BP:


"BP=2x+140+x=140+3x."

Also, we have


"BC=\frac{CD}{2}=140+x."

And we also notice that


"BP=BC+PC=(140+x)+2x=140+3x."